An object is moving with a uniform acceleration a= 2ms^-2, determine the functions for:
A) velocity v (ms^-1) - given v (0) = 10ms^-1 ?
B) displacement s (m) - given s (0) = 5m ?
C) calculate the values of v and s for:
i) t = 2s
ii) t = 5s
1
Expert's answer
2019-11-26T14:07:12-0500

The equations of motion with constant acceleration are:

"v(t) = v_0 + a t", "s(t) = s_0 + v_0 t + \frac{a t^2}{2}"

A) Using the first equation, the velocity function is "v(t) = 10 + 2 t".

B) Using the second equation, displacement function is "s(t) = 5 +10 t + t^2".

C) "v(2) = 10 \frac{m}{s} + 2 \frac{m}{s^2}\cdot 2 s = 14 \frac{m}{s}" , "s(2) = 5 m + 10\frac{m}{s} \cdot 2 s + 1 \frac{m}{s^2} \cdot 2^2 s^2 = 29 m".

D) "v(5) = 10\frac{m}{s} + 2\frac{m}{s^2} \cdot 5 s = 20 \frac{m}{s}" , "s(5) = 5 m + 10\frac{m}{s} \cdot 5 s + 1\frac{m}{s^2} \cdot 5^2 s^2 = 80 m".